# Reaction Prediction via Interaction Modeling of Symmetric Difference Shingle Sets

### v2

@@ 1 Introduction
Accurate modeling of chemical reactions is a fundamental problem in organic chemistry, as it provides critical insights into reaction mechanisms, predicts reaction outcomes, and guides experimental design [organic_0, organic_1, organic_2, organic_3]. Reaction representation learning is central to tasks such as reaction yield prediction [rea_bert], enantioselectivity prediction [selectivity_0], conversion rate estimation [rate_0], and reaction type classification [mapping]. These tasks have gained considerable attention with the rise of machine learning (ML). Nevertheless, representing reactions for ML is challenging due to the complexity of reaction spaces and the multitude of factors influencing chemical experiments like substrates, catalysts, and reaction conditions [rea_hard_0, rea_hard_1, Yang2025BatGPT]. While many reactions may appear theoretically feasible [importance], in practice, successfully executing them requires a deeper understanding of how these factors interact. Even slight variations in any of these elements can significantly influence the outcome [rea_BH, rea_SM], making the reaction prediction problem a complex and nuanced challenge.
@@
Despite advances in ML models, they still face two major limitations that hinder their broader applicability, as shown in Figure 1. First, many models, especially those based on sequential representations like SMILES [smiles], fail to account for the inherent permutation invariance of chemical reactions [permutation_2, permutation_3]. They tend to produce inconsistent predictions when changing the ordering of input molecules (e.g., swapping reactants and reagents) or the ordering of atoms (e.g., alternative SMILES). Such sensitivity to input ordering undermines reliability and generalizability. While data augmentation techniques can partially alleviate this issue, training on all possible permutations is computationally inefficient and impractical at scale [permutation_0, rea_tf]. Achieving permutation invariance is essential for robust and unbiased reaction modeling across diverse chemical scenarios.
@@
Second, current works often fail to capture the structural interactions that drive chemical reactivity in reaction modeling. Many approaches overlook critical substructures, particularly those that change from reactants to products (e.g., reaction centers [center]), and lack explicit modeling of interactions between such substructures. Although atom-level interactions have been widely explored in molecular representation learning [unimol, pair_0, pair_1], capturing higher-order interactions among chemically meaningful substructures across molecules remains limited in the reaction prediction problem. This lack of interaction awareness leads to suboptimal predictions and diminishes the model’s ability to capture underlying reaction mechanisms [transformation]. As a result of these two limitations, many models perform poorly on out-of-sample reactions that better reflect real-world chemical diversity [BH_test_fail, SM_test_fail, challenge].
@@
To address these challenges, we propose the ReaDISH model for Reaction prediction via interaction modeling of symmetric DIfference SHingle sets. First, we construct chemically meaningful substructure sets named molecular shingle [shingle_0, shingle_1, shingle_2] sets via the circular topology for each molecule to capture key structural components. Then we compute and encode the symmetric difference of shingle sets between reactants and products adopted from DRFP [rea_DRFP], which is naturally robustly permutation-invariant. Second, we add intra- and inter-molecular interaction pair representation on shingle-level attention based on geometric distance, structural distance, and chemical connectivity of shingles to improve interaction awareness. We conduct numerical experiments on various reaction representation learning tasks and benchmarks, including reaction yield prediction, enantioselectivity prediction, conversion rate estimation, and reaction type classification. Experimental results show that compared with baseline models, our proposed ReaDISH achieves SOTA on prediction accuracy and uncertainty estimation in most scenarios. It increases R2 by an average of 8.76% under out-of-sample splits when performing permutation perturbations, presenting a better generalization capability. We present related work in Appendix A. Our key contributions are summarized as follows:
@@ 2.2 Molecular shingles
In cheminformatics, molecular shingles refer to structured fragments designed to capture local connectivity patterns in molecules. These shingles are typically defined as sets of atoms and their connectivity within a defined neighborhood, which are also utilized in classical fingerprints like ECFP [ECFP], effectively representing the structural information of the molecule.
@@ 3.1 Model architecture
Given a reaction $\mathcal{R}=\left\{C_{i}^{\text{r}_{1}},\dots,C_{i}^{\text{r}_{m}};C_{i}^{\text{p}_{1}},\dots,C_{i}^{\text{p}_{n}}\right\}$ consisting of reactant and product molecules in 3D conformer format, ReaDISH first encodes each molecule $C_{i}^{j}$ using a 3D molecular encoder $f_{\text{mol}}$ to generate atom-level representations $\mathbf{X}^{\text{a}}\in\mathbb{R}^{N_{\text{a}}\times F}$, where $F$ is the embedding dimension. A shingle-generation algorithm then extracts molecular shingles and computes the symmetric difference between reactant and product shingle sets to capture reaction-specific transformations. Next, ReaDISH aggregates the atom-level representations within each shingle through a pooling operation, producing initial shingle-level embeddings $\mathbf{X}^{0}\in\mathbb{R}^{N_{\text{s}}\times F}$, where $N_{\text{s}}$ is the number of shingles. We compute three types of pairwise relations to model interactions between shingles: geometric distances, structural distances, and chemical connectivity. These are encoded as pairwise matrices $\mathbf{P}=\{\mathbf{P}_{\text{g}},\mathbf{P}_{\text{e}},\mathbf{P}_{\text{s}}\}\in\mathbb{R}^{3\times N_{\text{s}}\times N_{\text{s}}}$. To incorporate these relationships into the attention mechanism, we apply $K$ Gaussian kernels to each pairwise matrix to produce the initial pair representation $\mathbf{P}^{0}$. The ReaDISH encoder then processes the shingle-level representations using $L$ transformer-style [transformer] attention blocks. Each block contains a multi-head attention (MHA) layer that integrates the pairwise attention biases $\mathbf{P}^{\ell}$ at each layer $\ell$. Finally, ReaDISH uses a special [SSUM] token to summarize the reaction representation and passes it to a multilayer perceptron (MLP) head to predict reaction property $\hat{y}$.
@@ 3.2 Embedding layer for shingles
In reaction modeling, comparing shingles across reactants and products allows for the precise identification of structural transformations [rea_DRFP]. For each reaction, we treat molecules except products as reactants. Shingles within a given radius are generated for reactants and products, respectively. The generated shingles are expressed in SMILES strings to perform the symmetric difference operation, as shown in Figure 3. This operation yields the set of shingles that are exclusive to either reactants or products, thereby identifying structural changes associated with the reaction mechanism. The complete algorithm and illustration for computing the symmetric difference shingle set are provided in Appendix C. Formally, for a reaction $\mathcal{R}=\{\mathcal{M}^{\text{r}};\mathcal{M}^{\text{p}}\}$ and a radius $r_{\text{max}}$, we define the symmetric difference shingle set $\mathcal{S}_{\mathcal{R}}$ as
@@
where $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ denote the geometric centers of shingles $S_{i}$ and $S_{j}$, respectively, and $\text{sim}(\cdot,\cdot)$ is the Tanimoto similarity [Tanimoto] of their Morgan fingerprints [Morgan]. These metrics yield a structured pairwise representation $\mathbf{P}=\{\mathbf{P}_{\text{g}},\mathbf{P}_{\text{e}},\mathbf{P}_{\text{s}}\}=\left[\,d_{ij}^{\text{g}},d_{ij}^{\text{e}},d_{ij}^{\text{s}}\,\right]_{i,j=1}^{N_{\text{s}}}\in\mathbb{R}^{3\times N_{\text{s}}\times N_{\text{s}}}$, where each component captures a specific interaction modality across all $N$ shingles. The geometric distance $d_{ij}^{\text{g}}$ encodes spatial relationships within the same molecule, analogous to atom-level distance modeling. The binary chemical connectivity $d_{ij}^{\text{e}}$ indicates whether two shingles are part of the same molecule. The structural distance $d_{ij}^{\text{s}}$ emphasizes functional dissimilarity, which is essential for modeling reactive interactions across different molecules. These pair interactions enable a finer-grained understanding of reaction mechanisms and facilitate more accurate prediction of reaction outcomes.
@@
After computing the pairwise interactions, we integrate these geometry-aware and structure-aware signals into the attention mechanism to better capture chemically meaningful context. To this end, we adopt the Gaussian Kernel with Pair Type (GKPT) [kernel], a technique effectively applied in molecular representation learning [unimol]. GKPT applies an affine transformation to pairwise distances based on the interaction type, followed by a classical Gaussian kernel, as illustrated in Figure 4(a). Formally, the GKPT is defined as
@@ 3.4 Predictor
ReaDISH utilizes a lightweight predictor, as depicted in Figure 2(d), to predict reaction properties. It consists of two linear layers with GELU activation [GELU]. This simple yet effective architecture maps the final [SSUM] token representation, which encodes the global context of the reaction, to the target prediction space.
@@ 4.1 Settings
We collect 3.7M chemical reactions for pre-training based on the United States Patent and Trademark Office (USPTO) dataset [USPTO] and the Chemical Journals with High Impact Factor (CJHIF) dataset [CJHIF]. We employ the DRFP [rea_DRFP] method with the $K$-means algorithm to compute pseudo-labels. More information can be found in Appendix B.
@@
To comprehensively evaluate the effectiveness of ReaDISH, we use seven datasets across a wide range of chemical tasks, including: (1) yield prediction, the Buchwald-Hartwig (BH) dataset [rea_BH], the Suzuki-Miyaura (SM) dataset [rea_SM], the real-world electronic laboratory notebook (ELN) dataset [rea_gnn_desc], and the Ni-catalyzed C-O bond activation (NiCOlit) dataset [rea_NiCOlit]; (2) enantioselectivitiy prediction, the asymmetric N,S-acetal formation (N,S-acetal) dataset [rea_NS]; (3) conversion rate estimation, the C-heteroatom-coupling reactions (C-heteroatom) dataset [rea_heteroatom]; and (4) reaction type classification, the USPTO_TPL dataset [mapping]. We consider both random and more challenging out-of-sample splits, which better reflect practical deployment scenarios. More information can be found in Appendix B.
@@
We use five representative models as baselines in our experiments: YieldBERT [rea_bert], YieldBERT-DA [rea_bert_augmentation], UA-GNN [rea_gnn_sum], UAM [rea_uam], and ReaMVP [rea_mvp], which we introduce in related work in Appendix A. We additionally compare our method with a cross-attention baseline that directly models reactant-product interactions as provided in Appendix E. Given the importance of spatial structural patterns in molecules, we adopt the SE(3)-invariant [se3] Uni-Mol [unimol] as our 3D molecular encoder $f_{\text{mol}}$ for its excellent performance. For regression tasks, we report mean absolute error (MAE $\downarrow$), root mean squared error (RMSE $\downarrow$), and coefficient of determination (R2 $\uparrow$). For classification tasks, we measure accuracy (ACC $\uparrow$), Matthews correlation coefficient (MCC $\uparrow$), and confusion entropy (CEN $\downarrow$). More information can be found in Appendix F.
@@ 4.2 Main results
On the particularly challenging ELN dataset, however, all methods show limited performance with R2 below 0.3. This underscores the difficulty in generalizing to complex, real-world reaction data. The poor results can be attributed to two main factors: (1) the ELN dataset includes a broader range of substrates, ligands, and solvents than other datasets, with significantly greater structural variability, making it harder for models to capture meaningful patterns [SM_test_fail]; and (2) limited fine-tuning data, which constrains the model’s ability to adapt to the diverse chemical contexts. Performance regarding training size is presented in Appendix G. Full results are available in Appendix H.
@@
A key strength of ReaDISH is its permutation-invariant design, which ensures consistent predictions regardless of the ordering of input molecules or SMILES tokens. To test whether this property enhances robustness in out-of-sample settings, we measure the standard deviation of model predictions across five runs, each using different random molecule orderings and SMILES permutations, the same as YieldBERT-DA [rea_bert_augmentation]. As shown in Table 1, ReaDISH consistently shows the lowest prediction variance. It increases R2 by 11.11% and 6.40% on the BH and SM out-of-sample splits (average 8.76%), respectively. In contrast, models without permutation invariance show greater variability and reduced performance under input perturbations. We show an example in Figure 6.
@@ Appendix A Related work
Chemical reactions involve complex transformations among multiple molecular entities, such as reactants, catalysts, and reagents. Accurate prediction of reaction outcomes requires expressive and chemically informed representations. Early approaches rely heavily on handcrafted features as molecular fingerprint features derived from domain knowledge, to encode atomic and physicochemical properties [rea_BH, rea_SM, rea_MFF_fp, rea_gnn_desc, rea_NiCOlit]. Recent advances in deep learning for reaction prediction can be grouped into three main categories: sequence-based, graph-based, and conformer-based models. Sequence-based models represent chemical reactions using linear notations such as SMILES and apply neural sequence architectures to learn patterns in tokenized strings. Notable examples include YieldBERT [rea_bert] and YieldBERT-DA [rea_bert_augmentation], which leverage BERT-based [BERT] encoders; MolFormer [rea_tf], which adopts a transformer [transformer] architecture; and T5Chem [rea_t5], which builds on the T5 language model [t5]. Graph-based methods encode molecular structures as graphs and extract meaningful representations using graph neural networks (GNNs). Rxn Hypergraph [permutation_3] employs hypergraph attention to capture molecular interactions. UA-GNN [rea_gnn_sum] sums reactant embeddings via GNNs and concatenates them with product representations. UAM [rea_uam] combines multi-view inputs, including graphs, SMILES, and molecular fingerprints, and aggregates their embeddings. LocalTransform [rea_gnn_template] focuses on local structural changes by encoding atom-level differences between reactants and products. Conformer-based models incorporate 3D atomic coordinates to learn geometry-aware representations, which can better capture stereoelectronic factors influencing reactivity. ReaMVP [rea_mvp] combines Bi-GRU networks [gru] with multi-view pre-training on conformers and SMILES. YieldFCP [rea_yieldfcp] introduces a cross-modal projector to align conformer and SMILES representations for yield prediction. Despite their respective advantages, these methods often overlook the rich intra- and inter-molecular interactions that influence reaction outcomes. In this work, we address this gap by introducing a geometry- and structure-enhanced modeling approach that explicitly incorporates such interactions into the representation learning process.
@@
Substructures (functional groups, motifs, and fragments) play a pivotal role in determining molecular properties and, by extension, reaction outcomes. Several fragment-based approaches have been proposed to decompose molecules into chemically meaningful parts. RECAP [RECAP] introduces predefined cleavage rules to generate fragments at drug-like bond types, while BRICS [BRICS] focuses on retrosynthetically relevant splits. ReLMole [ReLMole] further generalizes this process using graph-based heuristics to automatically extract relevant substructures. In parallel, fingerprinting methods such as ECFP [ECFP] iteratively encode local atomic environments, capturing circular subgraphs of increasing radii. These subgraphs, referred to as molecular shingles [shingle_0, shingle_1, shingle_2], serve as a compact and effective representation of chemical structures. DRFP [rea_DRFP] extracts shingles and applies a symmetric difference operation on reactant and product shingle sets to obtain reaction fingerprints. Unlike DRFP, our method directly models the interaction-aware symmetric difference via a transformer, allowing it to capture fine-grained transformations at the shingle level.

### v3

@@ 1 Introduction
Accurate modeling of chemical reactions is a fundamental problem in organic chemistry, as it provides critical insights into reaction mechanisms, predicts reaction outcomes, and guides experimental design [11, 12, 28, 60]. Reaction representation learning is central to tasks such as reaction yield prediction [46], enantioselectivity prediction [37], conversion rate estimation [21], and reaction type classification [45]. These tasks have gained considerable attention with the rise of machine learning (ML). Nevertheless, representing reactions for ML is challenging due to the complexity of reaction spaces and the multitude of factors influencing chemical experiments like substrates, catalysts, and reaction conditions [53, 47, 61]. While many reactions may appear theoretically feasible [18], in practice, successfully executing them requires a deeper understanding of how these factors interact. Even slight variations in any of these elements can significantly influence the outcome [1, 33], making the reaction prediction problem a complex and nuanced challenge.
@@
Despite advances in ML models, they still face two major limitations that hinder their broader applicability, as shown in Figure 1. First, many models, especially those based on sequential representations like SMILES [59], fail to account for the inherent permutation invariance of chemical reactions [55, 54]. They tend to produce inconsistent predictions when changing the ordering of input molecules (e.g., swapping reactants and reagents) or the ordering of atoms (e.g., alternative SMILES). Such sensitivity to input ordering undermines reliability and generalizability. While data augmentation techniques can partially alleviate this issue, training on all possible permutations is computationally inefficient and impractical at scale [4, 44]. Achieving permutation invariance is essential for robust and unbiased reaction modeling across diverse chemical scenarios.
@@
Second, current works often fail to capture the structural interactions that drive chemical reactivity in reaction modeling. Many approaches overlook critical substructures, particularly those that change from reactants to products (e.g., reaction centers [57]), and lack explicit modeling of interactions between such substructures. Although atom-level interactions have been widely explored in molecular representation learning [63, 10, 58], capturing higher-order interactions among chemically meaningful substructures across molecules remains limited in the reaction prediction problem. This lack of interaction awareness leads to suboptimal predictions and diminishes the model’s ability to capture underlying reaction mechanisms [52]. As a result of these two limitations, many models perform poorly on out-of-sample reactions that better reflect real-world chemical diversity [15, 3, 5].
@@
To address these challenges, we propose the ReaDISH model for Reaction prediction via interaction modeling of symmetric DIfference SHingle sets. First, we construct chemically meaningful substructure sets named molecular shingle [34, 6, 30] sets via the circular topology for each molecule to capture key structural components. Then we compute and encode the symmetric difference of shingle sets between reactants and products adopted from DRFP [35], which is naturally robustly permutation-invariant. Second, we add intra- and inter-molecular interaction pair representation on shingle-level attention based on geometric distance, structural distance, and chemical connectivity of shingles to improve interaction awareness. We conduct numerical experiments on various reaction representation learning tasks and benchmarks, including reaction yield prediction, enantioselectivity prediction, conversion rate estimation, and reaction type classification. Experimental results show that compared with baseline models, our proposed ReaDISH achieves SOTA on prediction accuracy and uncertainty estimation in most scenarios. It increases R2 by an average of 8.76% under out-of-sample splits when performing permutation perturbations, presenting a better generalization capability. We present related work in Appendix A. Our key contributions are summarized as follows:
@@ 2.2 Molecular shingles
In cheminformatics, molecular shingles refer to structured fragments designed to capture local connectivity patterns in molecules. These shingles are typically defined as sets of atoms and their connectivity within a defined neighborhood, which are also utilized in classical fingerprints like ECFP [39], effectively representing the structural information of the molecule.
@@ 3.1 Model architecture
Given a reaction $\mathcal{R}=\left\{C_{i}^{\text{r}_{1}},\dots,C_{i}^{\text{r}_{m}};C_{i}^{\text{p}_{1}},\dots,C_{i}^{\text{p}_{n}}\right\}$ consisting of reactant and product molecules in 3D conformer format, ReaDISH first encodes each molecule $C_{i}^{j}$ using a 3D molecular encoder $f_{\text{mol}}$ to generate atom-level representations $\mathbf{X}^{\text{a}}\in\mathbb{R}^{N_{\text{a}}\times F}$, where $F$ is the embedding dimension. A shingle-generation algorithm then extracts molecular shingles and computes the symmetric difference between reactant and product shingle sets to capture reaction-specific transformations. Next, ReaDISH aggregates the atom-level representations within each shingle through a pooling operation, producing initial shingle-level embeddings $\mathbf{X}^{0}\in\mathbb{R}^{N_{\text{s}}\times F}$, where $N_{\text{s}}$ is the number of shingles. We compute three types of pairwise relations to model interactions between shingles: geometric distances, structural distances, and chemical connectivity. These are encoded as pairwise matrices $\mathbf{P}=\{\mathbf{P}_{\text{g}},\mathbf{P}_{\text{e}},\mathbf{P}_{\text{s}}\}\in\mathbb{R}^{3\times N_{\text{s}}\times N_{\text{s}}}$. To incorporate these relationships into the attention mechanism, we apply $K$ Gaussian kernels to each pairwise matrix to produce the initial pair representation $\mathbf{P}^{0}$. The ReaDISH encoder then processes the shingle-level representations using $L$ transformer-style [56] attention blocks. Each block contains a multi-head attention (MHA) layer that integrates the pairwise attention biases $\mathbf{P}^{\ell}$ at each layer $\ell$. Finally, ReaDISH uses a special [SSUM] token to summarize the reaction representation and passes it to a multilayer perceptron (MLP) head to predict reaction property $\hat{y}$.
@@ 3.2 Embedding layer for shingles
In reaction modeling, comparing shingles across reactants and products allows for the precise identification of structural transformations [35]. For each reaction, we treat molecules except products as reactants. Shingles within a given radius are generated for reactants and products, respectively. The generated shingles are expressed in SMILES strings to perform the symmetric difference operation, as shown in Figure 3. This operation yields the set of shingles that are exclusive to either reactants or products, thereby identifying structural changes associated with the reaction mechanism. The complete algorithm and illustration for computing the symmetric difference shingle set are provided in Appendix C. Formally, for a reaction $\mathcal{R}=\{\mathcal{M}^{\text{r}};\mathcal{M}^{\text{p}}\}$ and a radius $r_{\text{max}}$, we define the symmetric difference shingle set $\mathcal{S}_{\mathcal{R}}$ as
@@
where $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ denote the geometric centers of shingles $S_{i}$ and $S_{j}$, respectively, and $\text{sim}(\cdot,\cdot)$ is the Tanimoto similarity [2] of their Morgan fingerprints [29]. These metrics yield a structured pairwise representation $\mathbf{P}=\{\mathbf{P}_{\text{g}},\mathbf{P}_{\text{e}},\mathbf{P}_{\text{s}}\}=\left[\,d_{ij}^{\text{g}},d_{ij}^{\text{e}},d_{ij}^{\text{s}}\,\right]_{i,j=1}^{N_{\text{s}}}\in\mathbb{R}^{3\times N_{\text{s}}\times N_{\text{s}}}$, where each component captures a specific interaction modality across all $N$ shingles. The geometric distance $d_{ij}^{\text{g}}$ encodes spatial relationships within the same molecule, analogous to atom-level distance modeling. The binary chemical connectivity $d_{ij}^{\text{e}}$ indicates whether two shingles are part of the same molecule. The structural distance $d_{ij}^{\text{s}}$ emphasizes functional dissimilarity, which is essential for modeling reactive interactions across different molecules. These pair interactions enable a finer-grained understanding of reaction mechanisms and facilitate more accurate prediction of reaction outcomes.
@@
After computing the pairwise interactions, we integrate these geometry-aware and structure-aware signals into the attention mechanism to better capture chemically meaningful context. To this end, we adopt the Gaussian Kernel with Pair Type (GKPT) [51], a technique effectively applied in molecular representation learning [63]. GKPT applies an affine transformation to pairwise distances based on the interaction type, followed by a classical Gaussian kernel, as illustrated in Figure 4(a). Formally, the GKPT is defined as
@@ 3.4 Predictor
ReaDISH utilizes a lightweight predictor, as depicted in Figure 2(d), to predict reaction properties. It consists of two linear layers with GELU activation [17]. This simple yet effective architecture maps the final [SSUM] token representation, which encodes the global context of the reaction, to the target prediction space.
@@ 4.1 Settings
We collect 3.7M chemical reactions for pre-training based on the United States Patent and Trademark Office (USPTO) dataset [26] and the Chemical Journals with High Impact Factor (CJHIF) dataset [20]. We employ the DRFP [35] method with the $K$-means algorithm to compute pseudo-labels. More information can be found in Appendix B.
@@
To comprehensively evaluate the effectiveness of ReaDISH, we use seven datasets across a wide range of chemical tasks, including: (1) yield prediction, the Buchwald-Hartwig (BH) dataset [1], the Suzuki-Miyaura (SM) dataset [33], the real-world electronic laboratory notebook (ELN) dataset [40], and the Ni-catalyzed C-O bond activation (NiCOlit) dataset [43]; (2) enantioselectivitiy prediction, the asymmetric N,S-acetal formation (N,S-acetal) dataset [62]; (3) conversion rate estimation, the C-heteroatom-coupling reactions (C-heteroatom) dataset [42]; and (4) reaction type classification, the USPTO_TPL dataset [45]. We consider both random and more challenging out-of-sample splits, which better reflect practical deployment scenarios. More information can be found in Appendix B.
@@
We use five representative models as baselines in our experiments: YieldBERT [46], YieldBERT-DA [48], UA-GNN [23], UAM [7], and ReaMVP [50], which we introduce in related work in Appendix A. We additionally compare our method with a cross-attention baseline that directly models reactant-product interactions as provided in Appendix E. Given the importance of spatial structural patterns in molecules, we adopt the SE(3)-invariant [16] Uni-Mol [63] as our 3D molecular encoder $f_{\text{mol}}$ for its excellent performance. For regression tasks, we report mean absolute error (MAE $\downarrow$), root mean squared error (RMSE $\downarrow$), and coefficient of determination (R2 $\uparrow$). For classification tasks, we measure accuracy (ACC $\uparrow$), Matthews correlation coefficient (MCC $\uparrow$), and confusion entropy (CEN $\downarrow$). More information can be found in Appendix F.
@@ 4.2 Main results
On the particularly challenging ELN dataset, however, all methods show limited performance with R2 below 0.3. This underscores the difficulty in generalizing to complex, real-world reaction data. The poor results can be attributed to two main factors: (1) the ELN dataset includes a broader range of substrates, ligands, and solvents than other datasets, with significantly greater structural variability, making it harder for models to capture meaningful patterns [3]; and (2) limited fine-tuning data, which constrains the model’s ability to adapt to the diverse chemical contexts. Performance regarding training size is presented in Appendix G. Full results are available in Appendix H.
@@
A key strength of ReaDISH is its permutation-invariant design, which ensures consistent predictions regardless of the ordering of input molecules or SMILES tokens. To test whether this property enhances robustness in out-of-sample settings, we measure the standard deviation of model predictions across five runs, each using different random molecule orderings and SMILES permutations, the same as YieldBERT-DA [48]. As shown in Table 1, ReaDISH consistently shows the lowest prediction variance. It increases R2 by 11.11% and 6.40% on the BH and SM out-of-sample splits (average 8.76%), respectively. In contrast, models without permutation invariance show greater variability and reduced performance under input perturbations. We show an example in Figure 6.
@@ Appendix A Related work
Chemical reactions involve complex transformations among multiple molecular entities, such as reactants, catalysts, and reagents. Accurate prediction of reaction outcomes requires expressive and chemically informed representations. Early approaches rely heavily on handcrafted features as molecular fingerprint features derived from domain knowledge, to encode atomic and physicochemical properties [1, 33, 41, 40, 43]. Recent advances in deep learning for reaction prediction can be grouped into three main categories: sequence-based, graph-based, and conformer-based models. Sequence-based models represent chemical reactions using linear notations such as SMILES and apply neural sequence architectures to learn patterns in tokenized strings. Notable examples include YieldBERT [46] and YieldBERT-DA [48], which leverage BERT-based [14] encoders; MolFormer [44], which adopts a transformer [56] architecture; and T5Chem [27], which builds on the T5 language model [36]. Graph-based methods encode molecular structures as graphs and extract meaningful representations using graph neural networks (GNNs). Rxn Hypergraph [54] employs hypergraph attention to capture molecular interactions. UA-GNN [23] sums reactant embeddings via GNNs and concatenates them with product representations. UAM [7] combines multi-view inputs, including graphs, SMILES, and molecular fingerprints, and aggregates their embeddings. LocalTransform [8] focuses on local structural changes by encoding atom-level differences between reactants and products. Conformer-based models incorporate 3D atomic coordinates to learn geometry-aware representations, which can better capture stereoelectronic factors influencing reactivity. ReaMVP [50] combines Bi-GRU networks [9] with multi-view pre-training on conformers and SMILES. YieldFCP [49] introduces a cross-modal projector to align conformer and SMILES representations for yield prediction. Despite their respective advantages, these methods often overlook the rich intra- and inter-molecular interactions that influence reaction outcomes. In this work, we address this gap by introducing a geometry- and structure-enhanced modeling approach that explicitly incorporates such interactions into the representation learning process.
@@
Substructures (functional groups, motifs, and fragments) play a pivotal role in determining molecular properties and, by extension, reaction outcomes. Several fragment-based approaches have been proposed to decompose molecules into chemically meaningful parts. RECAP [25] introduces predefined cleavage rules to generate fragments at drug-like bond types, while BRICS [13] focuses on retrosynthetically relevant splits. ReLMole [19] further generalizes this process using graph-based heuristics to automatically extract relevant substructures. In parallel, fingerprinting methods such as ECFP [39] iteratively encode local atomic environments, capturing circular subgraphs of increasing radii. These subgraphs, referred to as molecular shingles [34, 6, 30], serve as a compact and effective representation of chemical structures. DRFP [35] extracts shingles and applies a symmetric difference operation on reactant and product shingle sets to obtain reaction fingerprints. Unlike DRFP, our method directly models the interaction-aware symmetric difference via a transformer, allowing it to capture fine-grained transformations at the shingle level.